the length of the chord of the parabola y^2=x which is bisected at the point (2,1) is

Grade:12

1 Answers

Rinkoo Gupta
9 years ago

T = S1


y-0.5(x+2) = 1^2 - 2 = -1



2y - x - 2 = -2


or, x = 2y



Put in the equation,



y^2 = 2y


or, y = 0 , 2


x = 0, 4


Points where the line cuts the parabola = (0,0), (4,2)



Length = sqrt(4^2 + 2^2) = sqrt(16+4) = sqrt(20) = 2*sqrt(5)

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty